OpenStudy (anonymous):

How do i solve 3x^2+2x+7=0 ? Will give medal to whoever helps Please be patient with me kind of tired so my brain is not working well right now

OpenStudy (anonymous):

So we divide both sides by 7?

OpenStudy (anonymous):

Ohh okay thanks

OpenStudy (anonymous):

I hate these problems >.< but i'm good at them. It's stupid. Lol let me start over I lost my trin of thought, I know a better way, sorry /)_-

OpenStudy (anonymous):

Lol okay (:

OpenStudy (anonymous):

Reorder the terms: 7 + 2x + 3x2 = 0 Divide all terms by, 3 the coefficient of the squared term: Divide each side by '3'. 2.333333333 + 0.6666666667x + x2 = 0 Move the constant term to the right: Add '-2.333333333' to each side of the equation. 2.333333333 + 0.6666666667x + -2.333333333 + x2 = 0 + -2.333333333 Reorder the terms: 2.333333333 + -2.333333333 + 0.6666666667x + x2 = 0 + -2.333333333 Combine like terms: 2.333333333 + -2.333333333 = 0.000000000 0.000000000 + 0.6666666667x + x2 = 0 + -2.333333333 0.6666666667x + x2 = 0 + -2.333333333 Combine like terms: 0 + -2.333333333 = -2.333333333 0.6666666667x + x2 = -2.333333333 The x term is 0.6666666667x. Take half its coefficient (0.3333333334). Square it (0.1111111112) and add it to both sides. Add '0.1111111112' to each side of the equation. 0.6666666667x + 0.1111111112 + x2 = -2.333333333 + 0.1111111112 Reorder the terms: 0.1111111112 + 0.6666666667x + x2 = -2.333333333 + 0.1111111112 Combine like terms: -2.333333333 + 0.1111111112 = -2.2222222218 0.1111111112 + 0.6666666667x + x2 = -2.2222222218 Factor a perfect square on the left side: (x + 0.3333333334)(x + 0.3333333334) = -2.2222222218

OpenStudy (anonymous):

I don't like fractions... decimals are better to use. But 0.66667 = 2/3

OpenStudy (anonymous):

I gtg, sorry :/ I'd explain it but I have to go do my chores D:

OpenStudy (anonymous):

its okay thank you! (: bye

OpenStudy (anonymous):

I hope this even helped. This is how my teacher taught me to do these. She's kind of annoying.

OpenStudy (anonymous):

It did and lol yeah my math teacher is really annoying and she doesnt explain well :P

OpenStudy (anonymous):

Ok, What You Have There, The Equation, Is In Standard Form ax^2+bx+c=0 \[3x^2+2x+7=0\] Do You See The Similarity? ☺

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, Good. Now What You Want To Do With That Equation Is Plug It Into "The Quadratic Formula", Which Is... \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Do You Know What To Do Next?

OpenStudy (anonymous):

not really i learned this but i didnt understand it well.

OpenStudy (anonymous):

Ok Well See Look At This Equation Closely, \[3x^2+2x+7=0\] \[ax^2+bx+c=0\] Where 3=a 2=b 7=c Now All You Have To Do Is Plug In Those Numbers Into The Quadratic Formula. \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Do You Know What I Mean? :o

OpenStudy (anonymous):

yes that part I get, what I mean is how to solve it.

OpenStudy (anonymous):

Ok That Is Totally Fine. \[x=\frac{ -2\pm \sqrt{2^2-4\times3\times7} }{ 2\times3 }\] \[x=\frac{ -2\pm \sqrt{4-4\times3\times7} }{ 6 }\] \[x=\frac{ -2\pm \sqrt{4-84} }{ 6 }\] \[x=\frac{ -2\pm \sqrt{-80} }{ 6 }\]

OpenStudy (anonymous):

Are You Getting What To Do Now...?

OpenStudy (anonymous):

um i guess what is the + with the underline?

OpenStudy (anonymous):

Oh Ok, Well That Means Plus Or Minus. Which Means There Are Going To Be 2 Answers. First You Solve That Equation By Adding, Then Once You Get The Answer, You Do It Again But This Time You Subtract Instead.

OpenStudy (anonymous):

does the 80 stay a negative? cause im trying to find the square root of it and it keeps saying invalid input